OpenStudy (anonymous):

geometric mean of 3 and 48

OpenStudy (thejax):

To get the geometric mean of two numbers, multiply them and take the square root: √(3 * 48) = √144 = 12 Now I think it wants you to find the two geometric means of 3 & 12, and then 12 & 48: √(3 * 12) = √36 = 6 √(12 * 48) = √576 = 24 Answer: 3, 6, 12, 24, 48